A uniform diving board of mass 100kg and length 4m is supported at its left end and its center by two seperate supports. A 50kg child walks to the right end of the board and stands there. Assuming that the board does not bend or break find the necessary support forces on the left end support and the center support.
Answer #1

translational equilibrium

T2= tension in the left-hand side support, T1 is the tension in center support

T2 + T1 = 150 x 9.8

T2 + T1= 1470 N

rotattional equilbrium from left hand ,

100 (9.8) ( 2) + 50 (9.8) (4) = T 1( 2)

200 x 2 x 9.8 = 2T1

T1 = 1960 N

T2= -490 N